# ul. Pasteura 5, 02-093 Warsaw, Poland

### v1

@@ 5.2 Short Volume Lifetime (SVL-LVS)
Let us now consider the case where the tension varies. We first provide a simple, unrealistic scenario where a background inspired in volume-modulus kination is followed by the hot Big Bang. We obtain analytic expressions for the frequencies where the spectral index changes and for the maximum amplitude. We then proceed to study the more realistic case where we solve numerically the dynamical system in Eqs. (33). We can obtain an expression for the amplitude in terms of the number of e-folds of kination, tracker and matter domination. We find the interesting result that the predicted signal is larger for a larger hierarchy between initial and late vacuum, because it allows for a longer period of kination with varying tension. The standard LVS (even with decays given by Eq. (45)) does not allow for a large signal but other scenarios do.
@@ 5.2.1 Toy example
We now turn to a simple example which illustrates the basic features of the above analysis. Consider a network formed at very early times (say, $\Lambda=10^{-4}M_{P}$) in a background inspired by volume modulus kination (i.e: $G\mu\sim 1/t$ and $a\sim t^{1/3}$). In this simple example we assume that the period gives rise to standard radiation domination with constant tension, which is too strong an assumption to describe the scenarios of Sec 3. In addition, we operate under the unphysical assumption that the network is already in a scaling regime. Since, as we will see, the largest contribution to the spectrum arises from the earliest times, it would be interesting to compute the effects of the network formation or, if formed during inflation, the post-inflationary approach of the network to scaling Guedes:2018afo ; Cui:2019kkd ; Gouttenoire:2019kij ; Ferrer:2023uwz ; Datta:2025yow .
Figure 4: GW spectrum in an unrealistic scenario of volume modulus kination followed by the hot Big Bang, for different dependencies on the probability discussed in the literature, see the discussion around Eq. (17). Generically, $P<1$ boosts the spectrum. In addition, a dependence in $P$ of the initial size of the loops (encoded in $\chi$) can affect the amplitude and the frequencies of features in the spectrum. Experimental limits are not showed in the plot because the configuration is unrealistic.
@@
  • First turn. The transition between the standard, flat spectrum (associated to loops created and sourcing during radiation domination) to the steep one (associated to loops created and sourcing during volume modulus kination) can be computed by comparing the relative energy densities. Evaluating Eq. (73) in the radiation period ($m=n=4$) and Eq. (88) in the kination period ($m=n=6$, $p=q=1$) [^36] It is easy to show that the same spectral index is predicted for loops created during kination but sourcing during radiation ($m=6,\,n=4,\,p=1,\,q=0$) but the contribution is smaller., we find that the amplitudes are equal (in order of magnitude) at
    $f=f_{\Delta}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu_{\Delta}}\right)^{3/8}\,,$ (98)
@@
    $f_{peak}=f_{F}=\frac{2}{\alpha P^{\chi}}\frac{1}{t_{F}z_{F}}\,.$ (99)
@@
Before concluding our toy example, we note that we can compute the maximum amplitude in terms of the parameters by evaluating Eq. (88) at the peak. In the particular case of volume-modulus kination:
$h^{2}\Omega_{\rm{GW}}^{(\rm{peak})}\simeq 2A_{\Delta}\left(\frac{f_{F}}{f_{\Delta}}\right)^{4}\log[t_{\Delta}/t_{F}]\,.$ (100)
Figure 4 shows the spectrum in the toy model predicted by three different scenarios for probability-dependence. For $P<1$, the amplitude is boosted, although the factor depends on the equation of state of the background. For instance, the radiation plateau can be shown to depend on $A_{\Delta}P^{3\chi/2}$ (as can be easily shown by integrating Eq. (69) with integration limit given by Eq. (72)). On the other hand, the peak amplitude only depends on $P$ through $A_{\Delta}\sim P^{2(1-\beta)+\chi}$, which in our two benchmarks ($\beta=1/2$, $\chi=0$) and ($\beta=2/3$, $\chi=1/3$) give the same power of $P$, thus explaining that the peak amplitude is the same in these cases. The frequencies are however different, and this can be explained by Eqs. (98) and (99). According to Eq. (98), the flat-kination transition depends very weakly on $P$ as $f\sim P^{-5\chi/8}$, which explains that the $P=1$ and $\chi=0$ cases occur at a similar frequency, while the transition in the $\chi=1/3$ case occurs at a slightly larger frequency. The peak frequency in Eq. (99), however, depends on $P^{\chi}$ which explains the order of magnitude difference between the peak positions in the different scenarios.
@@ 5.2.2 Realistic scenario
Equation (100) computes the value of the GW spectrum sourced by volume-modulus kination at its high-frequency peak. We notice, in particular, that the details of the posterior cosmology (including the other expected non-standard epochs such as trackers and matter domination) are included in the redshift factors. We can thus use these equations to determine whether the high-frequency spectrum is observationally interesting/viable from the details of a given scenario; an example is shown in Fig. 5, and the scenario is described in Appendix A. In terms of numbers of e-folds:
$h^{2}\Omega_{\rm{GW}}^{(\rm{peak})}\simeq 5\times 10^{2}\times\frac{C_{\rm{eff}}}{30.6}\times\frac{\Gamma}{50}\times\frac{0.1}{\alpha}\times\frac{4\times 10^{-3}}{P^{2(1-\beta)+\chi}}\times 3N_{kin}\,\times\left(G\mu_{0}\right)^{2}\times\exp\left({2N_{kin}-N_{m}}\right)\,.$ (101)
Figure 5: GW spectrum predicted by the SVL-LVS scenario as described in Sec. 5.2.2 and App. A. The flat plateau characteristic of radiation domination is changed to a dip in the spectrum due to the early matter domination. This is preceded by a period of volume modulus kination, which induces a large high-frequency boost in the signal, with amplitude as in Eq. (101). The plot shows two different probability dependences quoted in the literature, see the discussion around Eq. (17). We also show CMB-S4 projected bounds CMB-S4:2020lpa .
@@
  • $N_{m}$. The number of matter efolds can be approximated noting that matter domination starts approximately at $t\sim 1/m_{\Phi}$ and the volume mode decays at $t_{reh}\sim 1/\Gamma$. Hence,
    $N_{m}\simeq\frac{2}{3}\ln\left(\frac{m_{\Phi}}{\Gamma_{\Phi}}\right)\,.$ (102)
@@
    $N_{m}\simeq 12+\frac{4}{3}\ln\left(\frac{10^{10}}{\mathcal{V}}\times\frac{30\text{ TeV}}{m_{\Phi}}\right)$ (103)
    whilst for gravitational decay we recall $\Gamma^{(g)}=\Gamma^{(f)}/\mathcal{V}^{2}$. In LARGE volume scenarios, this induces tens of additional efolds of matter domination, resulting in an unobservable signal.
  • $N_{kin}$. The number of e-folds of kination is somewhat harder to estimate. It is interesting, however, that the larger that the hierarchy between $V_{inf}$ and $V_{late}$ is, the larger is the field displacement and so the larger number of e-folds of kination. For illustrative purposes, inspired by Cicoli:2024bwq , we have considered a potential of the form $V=V_{0}e^{-\lambda\phi}$ with small $V_{0}=10^{-10}$ and $\lambda=4\sqrt{2/3}$ (see the Appendix A for details).
    The main obstruction to increasing $N_{kin}$ in this manner is that as $V_{late}$ gets smaller the easiest it is to overshoot, getting a phenomenologically unviable scenario. Thus this longer period where the fields falls down its potential must be compensated by additional efolds of tracker (which, as is evident from Eq. (101), do not affect the amplitude). We have observed that $N_{kin}$ grows logarithmically with the hierarchy $V_{inf}/V_{late}$, reaching around 4-5 efolds of kination in the scenario discussed in Appendix A but a more concrete analysis is left for future work.
We thus conclude that LARGE volume scenarios with hierarchically large separations between inflation (or other UV scale) and late-time scales featuring volume modulus kination source potentially interesting signals. Counterintuitively, it seems like a larger late-time volume (which leads to a longer period of early matter domination) yields a larger signal. This is due to several factors on top of the usual phenomenon of amplitudes getting larger because the background redshifts faster than the GWs (see e.g. Gouttenoire:2019kij ). Firstly, the early-time emission is much stronger due to the enhancement of the tension at early times. Second, the loop distribution is modified. Together, these facts render an approximately constant contribution to the spectrum at all relevant frequency bins during kination, which is manifest in the logarithmic behaviour in Eq. (100). This is different to the constant tension case where at every frequency bin of relevance the largest contribution arises at the time $t_{M}(f)$ for given $f$, as discussed in Sec. 4.1.
@@ 6 Discussion
Let us now conclude and summarize our main findings. After a review of cosmic strings in string theory and moduli dynamics in the early Universe in the context of LARGE volume scenarios, we have concluded that these models naturally accommodate the values of $G\mu$ and $P$ hinted at by PTA measurements. [^37] We have taken as a proxy the values quoted in Ellis:2023tsl as best-fit, which took into account LVK constraints. Our scenarios, however, predict a drop in the spectrum which invalidate these constraints. Studying the best-fit of Ellis:2023tsl without the LVK constraints remains for future work, although we can see from Eq. (97) that the drop in the spectrum will still lie in the LISA band for those values. At given $P$, larger values of the string tension would be excluded by PTA measurements, while smaller values would imply a volume modulus that is too light and hence in tension with BBN. Moreover, such models come with additional predictions for the GW spectrum across the whole frequency range. It is particularly interesting to observe how LVS scenarios where the volume mode decays gravitationally make a concrete prediction for a turning frequency which robustly lies within the LISA band, as shown in Figs. 2 and 3. Notice how, in this context, the epoch of early matter domination needed to explain a non-detection by LVK is a natural prediction of the stringy cosmology, and does need to be added by hand (unlike, for example, Ellis:2023tsl ).
@@
In order to provide quantitative estimates for the GW spectrum and its spectral index, we have generalized the approach in Cui:2017ufi to include the effects of varying tension, finding spectral indices that cannot be reproduced by cosmic strings with constant tension. In particular, we obtained analytic expressions for both spectral indices and peak frequency, Eqs. (88) and (89) respectively. These results are model-independent and can be used for any situation where the tension of the strings varies with time as a power law. Some possibilities have been discussed recently in Revello:2024gwa ; Brunelli:2025ems and applying our results to those scenarios is a natural future direction. Interestingly, for rapidly varying tension the GW spectrum at a given frequency is dominated by GW emission from the earliest times. This observation potentially provides a window into physics at very high energies, including the time of formation of the network.
@@
Lastly, taking the PTA-suggested values for the intercommutation probability and string tension, we have distinguished two well-motivated scenarios and discussed in detail the corresponding predictions. In the vanilla LVS scenario, we have argued that there is a prediction for a tilt in the GW spectrum in the LISA band due to a period of early matter domination where the volume mode dominates the energy density (see Figures 2 and 3). The prediction for the tilt frequency in scenarios featuring gravitational decay is given in Eq. (97) as a function of the parameter $\alpha$, the intercommutation probability and the mass of the volume mode. As discussed in the last paragraph of Sec. 5.1, the prediction is robust, even if the assumption that the PTA signal is explained by cosmic superstrings is lifted. In the event of a positive detection of a GW background from cosmic superstrings, the LISA mission could provide information about microscopic, stringy parameters of our own universe, including the size of the extra dimensions and the mass of its associated mode. [^38] Of course, this signal is degenerate with any other source of early matter domination ending slightly before BBN. However, having a hint of the mass of the volume mode and the size of the extra dimensions would suggest other complementary probes, such as collider signals.
@@
In addition, we have computed the peak amplitude of the spectrum sourced after a period with varying tension and matter domination in Eq. (101) (with peak frequency in Eq. (99)). The result is the same dependence in number of e-folds of kination and matter as would be expected in a case with constant tension, provided one evaluates the dimensionless tension $G\mu_{0}$ at the time of emission. The signal is thus potentially much larger (a ratio of $(G\mu_{0}/G\mu_{end})^{2}$) than in a case with constant tension. This motivated us to study the SVL-LVS case, where we have argued that LARGE volume scenarios with volume-modulus kination (and the consequent varying tension) in the early Universe potentially provide large signals provided there is a large hierarchy between the UV and late-time scales, as shown in Figure 5. This is another example of high-frequency GWs potentially providing access to the earliest times of the Universe, with larger signals sourced at earlier times Aggarwal:2020olq ; Aggarwal:2025noe . [^39] The paramount example is the Standard Model Ghiglieri:2015nfa ; Ghiglieri:2020mhm ; Muia:2023wru with a signal that grows with the reheating temperature, and the analogue in string theory Frey:2024jqy where the signal is larger than in the Standard Model and grows with the string scale. The main obstacles to claiming that this behavior is generic are the length of the matter-domination period and the overshoot problem. It would be very interesting to understand whether other scenarios, such as the ones discussed in Brunelli:2025ems , can improve the situation since the potentials and decay channels are different. We emphasize here that the loops source the largest contribution to the GW spectrum as they are formed, in contrast with the constant tension case.
@@
In the upcoming years Gravitational Wave astronomy aspires to achieve measurement precisions that will be orders of magnitude improved than the current GW detectors. It is conceivable that such measurements will also lead us closer to understanding fundamental aspects in the microscopic structure of our Universe. In particular, the signals we discuss in this article represent a way to test cosmological consequences of String Theory, and potentially probe the size and dynamics of the extra dimensions it predicts.

### v2

@@ 5.2 Short Volume Lifetime (SVL-LVS)
Let us now consider the case where the tension varies. We first provide a simple, unrealistic scenario where a background inspired in volume-modulus kination is followed by the hot Big Bang. We obtain analytic expressions for the frequencies where the spectral index changes and for the maximum amplitude. We then proceed to study the more realistic case where we solve numerically the dynamical system in Eqs. (3.12). We can obtain an expression for the amplitude in terms of the number of e-folds of kination, tracker and matter domination. We find the interesting result that the predicted signal is larger for a larger hierarchy between initial and late vacuum, because it allows for a longer period of kination with varying tension. The standard LVS (even with decays given by Eq. (3.24)) does not allow for a large signal but other scenarios do.
@@ 5.2.1 Toy example
We now turn to a simple example which illustrates the basic features of the above analysis. Consider a network formed at very early times (say, $\Lambda=10^{-4}M_{P}$) in a background inspired by volume modulus kination (i.e: $G\mu\sim 1/t$ and $a\sim t^{1/3}$). In this simple example we assume that the period gives rise to standard radiation domination with constant tension, which is too strong an assumption to describe the scenarios of Sec 3. In addition, we operate under the unphysical assumption that the network is already in a scaling regime. Since, as we will see, the largest contribution to the spectrum arises from the earliest times, it would be interesting to compute the effects of the network formation or, if formed during inflation, the post-inflationary approach of the network to scaling [184, 185, 30, 186, 187].
Figure 4: GW spectrum in an unrealistic scenario of volume modulus kination followed by the hot Big Bang, for different dependencies on the probability discussed in the literature, see the discussion around Eq. (2.17). Generically, $P<1$ boosts the spectrum. In addition, a dependence in $P$ of the initial size of the loops (encoded in $\chi$) can affect the amplitude and the frequencies of features in the spectrum. Experimental limits are not showed in the plot because the configuration is unrealistic.
@@
  • First turn. The transition between the standard, flat spectrum (associated to loops created and sourcing during radiation domination) to the steep one (associated to loops created and sourcing during volume modulus kination) can be computed by comparing the relative energy densities. Evaluating Eq. (4.15) in the radiation period ($m=n=4$) and Eq. (4.30) in the kination period ($m=n=6$, p=q=1p=q=1) [^40] It is easy to show that the same spectral index is predicted for loops created during kination but sourcing during radiation (m=6,n=4,p=1,q=0m=6,\,n=4,\,p=1,\,q=0) but the contribution is smaller., we find that the amplitudes are equal (in order of magnitude) at
    $f=f_{\Delta}\left(\frac{\alpha P^{\chi}}{\Gamma G\mu_{\Delta}}\right)^{3/8}\,$, (5.7)
@@
    $f_{peak}=f_{F}=\frac{2}{\alpha P^{\chi}}\frac{1}{t_{F}z_{F}}\,.$ (5.8)
@@
Before concluding our toy example, we note that we can compute the maximum amplitude in terms of the parameters by evaluating Eq. (4.30) at the peak. In the particular case of volume-modulus kination:
$h^{2}\Omega_{\rm{GW}}^{(\rm{peak})}\simeq 2A_{\Delta}\left(\frac{f_{F}}{f_{\Delta}}\right)^{4}\log[t_{\Delta}/t_{F}]\,.$ (5.9)
Figure 4 shows the spectrum in the toy model predicted by three different scenarios for probability-dependence. For $P<1$, the amplitude is boosted, although the factor depends on the equation of state of the background. For instance, the radiation plateau can be shown to depend on $A_{\Delta}P^{3\chi/2}$ (as can be easily shown by integrating Eq. (4.11) with integration limit given by Eq. (4.14)). On the other hand, the peak amplitude only depends on $P$ through $A_{\Delta}\sim P^{2(1-\beta)+\chi}$, which in our two benchmarks ($\beta=1/2$, $\chi=0$) and ($\beta=2/3$, $\chi=1/3$) give the same power of $P$, thus explaining that the peak amplitude is the same in these cases. The frequencies are however different, and this can be explained by Eqs. (5.7) and (5.8). According to Eq. (5.7), the flat-kination transition depends very weakly on $P$ as $f\sim P^{-5\chi/8}$, which explains that the $P=1$ and $\chi=0$ cases occur at a similar frequency, while the transition in the $\chi=1/3$ case occurs at a slightly larger frequency. The peak frequency in Eq. (5.8), however, depends on $P^{\chi}$ which explains the order of magnitude difference between the peak positions in the different scenarios.
@@ 5.2.2 Realistic scenario
Equation (5.9) computes the value of the GW spectrum sourced by volume-modulus kination at its high-frequency peak. We notice, in particular, that the details of the posterior cosmology (including the other expected non-standard epochs such as trackers and matter domination) are included in the redshift factors. We can thus use these equations to determine whether the high-frequency spectrum is observationally interesting/viable from the details of a given scenario; an example is shown in Fig. 5, for a choice of potential:
$10^{10}V=\left(C_{1}e^{-\sqrt{2/3}\frac{\Phi}{M_{P}}}-e^{-2\sqrt{2/3}\frac{\Phi}{M_{P}}}\right)^{2}\,,\qquad C_{1}^{3/2}=3\cdot 10^{-10}\,$, (5.10)
whose motivation is further described in the Appendix. The peak amplitude is, in terms of number of e-folds:
$h^{2}\Omega_{\rm{GW}}^{(\rm{peak})}\simeq 5\times 10^{2}\times\frac{C_{\rm{eff}}}{30.6}\times\frac{\Gamma}{50}\times\frac{0.1}{\alpha}\times\frac{4\times 10^{-3}}{P^{2(1-\beta)+\chi}}\times 3N_{kin}\,\times\left(G\mu_{0}\right)^{2}\times\exp\left({2N_{kin}-N_{m}}\right)\,.$ (5.11)
Figure 5: GW spectrum predicted by the SVL-LVS scenario as described in Sec. 5.2.2 and App. A. The flat plateau characteristic of radiation domination is changed to a dip in the spectrum due to the early matter domination. This is preceded by a period of volume modulus kination, which induces a large high-frequency boost in the signal, with amplitude as in Eq. (5.11). The plot shows two different probability dependencies quoted in the literature, see the discussion around Eq. (2.17). We also show CMB-S4 projected bounds [188].
@@
  • $N_{m}$. The number of matter e-folds can be approximated noting that matter domination starts approximately at $t\sim 1/m_{\Phi}$ and the volume mode decays at $t_{reh}\sim 1/\Gamma$. Hence,
    $N_{m}\simeq\frac{2}{3}\ln\left(\frac{m_{\Phi}}{\Gamma_{\Phi}}\right)\,.$ (5.12)
@@
    $N_{m}\simeq 12+\frac{4}{3}\ln\left(\frac{10^{10}}{\mathcal{V}}\times\frac{30\text{ TeV}}{m_{\Phi}}\right)$ (5.13)
    whilst for gravitational decay we recall $\Gamma^{(g)}=\Gamma^{(f)}/\mathcal{V}^{2}$. In LARGE volume scenarios, this induces tens of additional e-folds of matter domination, resulting in an unobservable signal.
  • $N_{kin}$. The number of e-folds of kination is somewhat harder to estimate. It is interesting, however, that the larger that the hierarchy between $V_{inf}$ and $V_{late}$ is, the larger is the field displacement and so the larger number of e-folds of kination. For illustrative purposes, inspired by [39], we have considered a potential of the form $V=V_{0}e^{-\lambda\phi}$ with small $V_{0}=10^{-10}$ and $\lambda=4\sqrt{2/3}$ (see the Appendix A for details).
    The main obstruction to increasing $N_{kin}$ in this manner is that as $V_{late}$ gets smaller the easiest it is to overshoot, getting a phenomenologically unviable scenario. Thus this longer period where the field falls down its potential must be compensated by additional e-folds of tracker (which, as is evident from Eq. (5.11), do not affect the amplitude). We have observed that $N_{kin}$ grows logarithmically with the hierarchy $V_{inf}/V_{late}$, reaching around 4-5 e-folds of kination in the scenario discussed in Appendix A but a more concrete analysis is left for future work.
We thus conclude that LARGE volume scenarios with hierarchically large separations between inflation (or other UV scale) and late-time scales featuring volume modulus kination source potentially interesting signals. Counterintuitively, it seems like a larger late-time volume (which leads to a longer period of early matter domination) yields a larger signal. This is due to several factors on top of the usual phenomenon of amplitudes getting larger because the background redshifts faster than the GWs (see e.g. [30]). Firstly, the early-time emission is much stronger due to the enhancement of the tension at early times. Second, the loop distribution is modified. Together, these facts render an approximately constant contribution to the spectrum at all relevant frequency bins during kination, which is manifest in the logarithmic behaviour in Eq. (5.9). This is different to the constant tension case where at every frequency bin of relevance the largest contribution arises at the time $t_{M}(f)$ for given $f$, as discussed in Sec. 4.1.
@@ 6 Discussion
Let us now conclude and summarize our main findings. After a review of cosmic strings in string theory and moduli dynamics in the early Universe in the context of LARGE volume scenarios, we have concluded that these models naturally accommodate the values of $G\mu$ and $P$ hinted at by PTA measurements. [^41] We have taken as a proxy the values quoted in [7] as best-fit, which took into account LVK constraints. Our scenarios, however, predict a drop in the spectrum which invalidate these constraints. Studying the best-fit of [7] without the LVK constraints remains for future work, although we can see from Eq. (5.6) that the drop in the spectrum will still lie in the LISA band for those values. At given $P$, larger values of the string tension would be excluded by PTA measurements, while smaller values would imply a volume modulus that is too light and hence in tension with BBN. Moreover, such models come with additional predictions for the GW spectrum across the whole frequency range. It is particularly interesting to observe how LVS scenarios where the volume mode decays gravitationally make a concrete prediction for a turning frequency which robustly lies within the LISA band, as shown in Figs. 2 and 3. Notice how, in this context, the epoch of early matter domination needed to explain a non-detection by LVK is a natural prediction of the stringy cosmology, and does need to be added by hand (unlike, for example, [7]).
@@
In order to provide quantitative estimates for the GW spectrum and its spectral index, we have generalized the approach in [189] to include the effects of varying tension, finding spectral indices that cannot be reproduced by cosmic strings with constant tension. In particular, we obtained analytic expressions for both spectral indices and peak frequency, Eqs. (4.30) and (4.31) respectively. These results are model-independent and can be used for any situation where the tension of the strings varies with time as a power law. Some possibilities have been discussed recently in [43, 80] and applying our results to those scenarios is a natural future direction. Interestingly, for rapidly varying tension the GW spectrum at a given frequency is dominated by GW emission from the earliest times. This observation potentially provides a window into physics at very high energies, including the time of formation of the network.
@@
Lastly, taking the PTA-suggested values for the intercommutation probability and string tension, we have distinguished two well-motivated scenarios and discussed in detail the corresponding predictions. In the vanilla LVS scenario, we have argued that there is a prediction for a tilt in the GW spectrum in the LISA band due to a period of early matter domination where the volume mode dominates the energy density (see Figures 2 and 3). The prediction for the tilt frequency in scenarios featuring gravitational decay is given in Eq. (5.6) as a function of the parameter $\alpha$, the intercommutation probability and the mass of the volume mode. As discussed in the last paragraph of Sec. 5.1, the prediction is robust, even if the assumption that the PTA signal is explained by cosmic superstrings is lifted. In the event of a positive detection of a GW background from cosmic superstrings, the LISA mission could provide information about microscopic, stringy parameters of our own universe, including the size of the extra dimensions and the mass of its associated mode. [^42] Of course, this signal is degenerate with any other source of early matter domination ending slightly before BBN. However, having a hint of the mass of the volume mode and the size of the extra dimensions would suggest other complementary probes, such as collider signals.
@@
In addition, we have computed the peak amplitude of the spectrum sourced after a period with varying tension and matter domination in Eq. (5.11) (with peak frequency in Eq. (5.8)). The result is the same dependence in number of e-folds of kination and matter as would be expected in a case with constant tension, provided one evaluates the dimensionless tension $G\mu_{0}$ at the time of emission. The signal is thus potentially much larger (a ratio of $(G\mu_{0}/G\mu_{end})^{2}$) than in a case with constant tension. This motivated us to study the SVL-LVS case, where we have argued that LARGE volume scenarios with volume-modulus kination (and the consequent varying tension) in the early Universe potentially provide large signals provided there is a large hierarchy between the UV and late-time scales, as shown in Figure 5. This is another example of high-frequency GWs potentially providing access to the earliest times of the Universe, with larger signals sourced at earlier times [40, 42]. [^43] The paramount example is the Standard Model [190, 191, 192] with a signal that grows with the reheating temperature, and the analogue in string theory [73] where the signal is larger than in the Standard Model and grows with the string scale. The main obstacles to claiming that this behavior is generic are the length of the matter-domination period and the overshoot problem. It would be very interesting to understand whether other scenarios, such as the ones discussed in [80], can improve the situation since the potentials and decay channels are different. We emphasize here that the loops source the largest contribution to the GW spectrum as they are formed, in contrast with the constant tension case.
@@
In the upcoming years Gravitational Wave astronomy aspires to achieve measurement precisions that will be orders of magnitude improved than the current GW detectors. It is conceivable that such measurements will also lead us closer to understanding fundamental aspects in the microscopic structure of our Universe. In particular, the signals we discuss in this article represent a way to test cosmological consequences of certain String Theory constructions, and potentially probe the size and dynamics of the extra dimensions it predicts.